### v2

# H-Fac: Memory-Efficient Optimization with Factorized Hamiltonian Descent

Abstract
In this study, we introduce a novel adaptive optimizer, H-Fac, which incorporates a factorized approach to momentum and scaling parameters. Our algorithm demonstrates competitive performances on both ResNets and Vision Transformers, while achieving sublinear memory costs through the use of rank-1 parameterizations for moment estimators. We develop our algorithms based on principles derived from Hamiltonian dynamics, providing robust theoretical underpinnings. These optimization algorithms are designed to be both straightforward and adaptable, facilitating easy implementation in diverse settings.
1 Introduction
Optimization algorithms play an indisputable role in the remarkable development of AI, especially in the realm of modern deep learning. In recent years, the emergence of breakthroughs in architectural innovation [3], as well as practical applications [37], has further promoted the necessity for embracing efficient training paradigms, which encompass optimization algorithms striking a balance between performance and manageable memory costs.
Stochastic gradient descent (SGD) is widely regarded as the standard algorithm for training deep learning models, supported by extensive theoretical foundations [31, 32, 34, 43]. However, it requires thorough tuning of hyperparameters and frequently exhibits undesirable convergence rates when applied to many contemporary architectures [10, 36, 40]. Meanwhile, adaptive gradient methods such as Adam [17], AdaGrad [12], AMSGrad [29], etc., can adjust the learning rate for each parameter throughout the optimization process by utilizing cumulative second-order statistics. Although the theoretical aspects have not yet been fully exploited, these methods show empirical performance that surpasses SGD across multiple domains and often provides better convergence properties in practice [41, 24], making them highly appealing for large-scale applications. However, maintaining momentum and per-coordinate scaling parameter accumulators in these algorithms will significantly increase memory overhead. This barrier typically restricts model size, reduces the number of examples per mini-batch, or limits communication in decentralized training [18, 21, 22], thereby negatively affecting convergence and accuracy.
Several memory-efficient optimizers have been devised to address this problem. In particular, Adafactor [30] is a highly efficient algorithm that was designed for low memory usage and greater scalability. It achieves substantial memory savings by employing the non-negative matrix factorization approach [19] to decompose the second-moment accumulator into two distinct rank-1 factors. This technique not only minimizes memory footprint but also enhances computational efficiency. However, to avoid performance degradation, Adafactor still needs to maintain first-order statistics, particularly when applied to large-scale models. This could potentially limit its applicability in memory-constrained environments. One of our goals is to address this inherent issue. Moreover, Adafactor does not yet have a solid theoretical foundation, which leaves its underlying methodologies and effectiveness without formal justification.
In this paper, we propose to approach the aforementioned open problems through the lens of Hamiltonian dynamics. Motivated by some inspirational works [4, 26], we explore how Hamiltonian principles in modeling dynamic systems can inform and improve optimization strategies, especially for memory efficiency purpose. Our contributions can be summarized in the following:
  • We demonstrate that Adafactor can be deduced from an ordinary differential equation (ODE) that solves a minimization problem with constrained factors. This novel perspective allows us to derive other algorithms with the same principles.
  • We propose a new class of efficient optimization algorithms that embraces both first-order and adaptive methods. By employing rank-1 parameterization for momentum and scaling parameter, our optimizers can offer sublinear memory costs, comparable to that of vanilla SGD without momentum. To the best of our knowledge, our proposal is the first endeavor to exploit Hamiltonian dynamics in developing a class of memory-efficient optimization algorithms that utilize factorized gradient statistic estimators.
  • Distinct from existing optimization techniques for memory efficiency, our algorithm can offer clear insights into optimization dynamics and convergence guarantees, which are naturally inherited from the fundamental theory of Hamiltonian mechanics.
  • Empirical results show that our optimization algorithms can achieve favorable and comparable results on ResNets and Vision Transformers.
Organization. The rest of the paper is organized as follows. The next section will discuss related works on memory-efficient optimization. We will provide in section 3 a brief overview of the Hamiltonian descent framework. Section 4 presents our main methodological contributions. The experimental results are shown in section 5. Finally, in section 6, we will discuss the current limitations, as well as the potential of our proposal for future works. The proofs and some other results are left in Appendix.
Notations: We denote model parameters by a matrix $X$ of size $m\times n$. Let $1_{m}$ and $1_{n}$ be vectors of ones with dimensions $m$ and $n$, respectively. For any matrices $X,Y$ of size $m\times n$, we use $\sqrt{X}$ for element-wise square root, $X^{2}$ for element-wise square, and $X/Y$ to denote element-wise division. $X^{\top}$ stands for the transpose matrix, $\text{trace}(X)$ denotes the trace of matrix $X$, and the function $\mathrm{RMS}(X)=\sqrt{\frac{1}{m\times n}\sum_{i,j}X^{2}_{ij}}$ represents root-mean-square calculation.
2 Related Works
In this section, we provide a brief overview of existing memory-efficient optimization methods.
Developing memory-efficient optimization algorithms is increasingly critical and attempts to minimize memory usage of optimizers are based on many different aspects. Adafactor [30] minimizes the memory cost to a sublinear level by factorizing second-order statistics using a row-column outer product. Luo et al. [25] introduced a confidence-guided strategy to mitigate erroneous updates and reduce instability in Adafactor training, while still retaining the same memory footprint. SM3 [1] is another efficient algorithm, which organizes the parameter space into sets and simplifies the maintenance of second-order statistics by computing the maximum squared gradient for each set.
There are several methods that focus on the low-rank structure of the gradient rather than the moment statistics [15, 42]. For instance, GaLore [42] periodically applies Singular Value Decomposition (SVD) to project the full gradients onto a lower-dimensional subspace, and subsequently utilize these projected gradients in adaptive optimization processes. Sketchy [13] leverages the Frequent Directions (FD) sketch technique to maintain a low-rank approximation of the gradient covariances matrix. Additionally, there is a line of work adapting quantization to reduce the memory cost of optimizer states [8, 9, 20].
For second-order (Hessian-based) methods, the memory constraints primarily arise from computing the (inverse) Hessian matrix. Some successful works have also resolved this bottleneck using Gauss-Newton decomposition [23] or Kronecker-based factorization [27].
3 Background
3.1 Hamiltonian Descent
We consider an unconstrained, continuous optimization problem $\min_{X\in\mathbb{R}^{m\times n}}f(X)$, with a proper differentiable and lower bounded objective $f:\mathbb{R}^{m\times n}\rightarrow\mathbb{R}$. In deep learning particularly, gradient-based optimization algorithms have been the de facto choice for solving such problems. Gradient descent with momentum (GDm) is one widely adopted such method due to its simplicity and effectiveness. Specifically, the update procedure of GDm is based on the following iterative scheme:
$M_{t}=\beta M_{t-1}+\nabla f(X_{t-1}),~{}~{}\quad X_{t}=X_{t-1}-\eta_{t}M_{t},$
where $\eta_{t}$ is the step size at the $t$-th iteration, $\beta\in(0,1)$ is a coefficient controlling how much we decay the old momentum at each step. In case we employ full-batch gradients, this scheme is deterministic and can be viewed as a discretization of a continuous-time system [26, 14] as follows
$\dot{M}=-\gamma M-\nabla f(X),~{}~{}\quad\dot{X}=M,$ (1)
where $M$ and $\gamma$ are in analogy to the velocity the friction in classical mechanics. In principle, this ordinary differential equation (ODE) defines a trajectory of the particle $X_{t}$ and its velocity $M_{t}$, which drives the systems characterized by the total energy or Hamiltonian function,
$H(X,M)=f(X)+\|M\|_{2}^{2}/2,$
towards stationary points. Indeed, we can show that the Hamiltonian $H_{t}$ is monotonically decreasing along the ODE trajectory:
$\frac{\mathrm{d}}{\mathrm{d}t}H=\text{trace}(\nabla^{\top}f(X)\dot{X})+\text{trace}(M^{\top}\dot{M})=-\gamma\text{trace}(M^{\top}M)=-\gamma\|M\|^{2}_{F}\leq 0.$
in which .$\left\lVert.\right\rVert_{F}$ is the Frobenius norm. This interpretation is meaningful in the sense that it facilitates us to theoretically establish a broader class of optimizers with guarantees, as well as providing complementary insights, improving or extending existing algorithms. In the next section, we will strengthen this argument by extending the Hamiltonian framework to more modern algorithms, and especially provide new memory-efficient optimizers based on this principle.
4 Factorized Hamiltonian Descent
In this section, we first interpret Adafactor optimizer from the perspective of Hamiltonian descent. We then introduce a general method to factorize the momentum in first-order optimization methods. We will specifically deliver a factorization version for the sign-based momentum update and name this optimizer as signFSGD. Based on further insights, we propose a novel adaptive factorized optimization method, named H-Fac. This is our main algorithm contribution to this paper.
4.1 Adafactor optimizer as Hamiltonian
Adafactor [30] shown in Algorithm 1 proposed an efficient rank-1 parameterization for the scaling factor, $V=rs^{\top}$, which is widely employed in adaptive optimization methods like Adam [17], RMSprop [35], etc. The updates of vectors $r$ and $s$ were inspired by the total elementwise I-divergence subject to componentwise non-negative constraints:
$\underset{r\in\mathbb{R}^{m},s\in\mathbb{R}^{n}}{\text{minimize}}\sum_{i=1}^{m}\sum_{j=1}^{n}d(V_{ij},r_{i}s_{j}),$
in which $r_{i}\geq 0,s_{j}\geq 0$ and $d(p,q)=p\log\frac{p}{q}-p+q$.
Algorithm 1 Adafactor-m for matrix parameter, with factored second moments and first-moment decay coefficient $\beta_{1}$.
  Inputs: moment decay coefficients $\beta_{1},\beta_{2}$, smoothing term $\epsilon$, and regularization constant $\lambda$
  Initialization: weight parameters $X_{0}\in\mathbb{R}^{m\times n}$, initial moments $M_{0},r_{0},s_{0}\leftarrow 0$
  for $t=1$ to $T$ do
     $G_{t}=\nabla f_{t}(X_{t-1})$
     $M_{t}=\hat{\beta}_{1t}M_{t-1}+(1-\hat{\beta}_{1t}G_{t}$
     $r_{t}=\hat{\beta}_{2t}r_{t-1}+(1-\hat{\beta}_{2t})\big{[}(G_{t})^{2}+\epsilon\big{]}1_{n}$
     $s_{t}=\hat{\beta}_{2t}s_{t-1}+(1-\hat{\beta}_{2t})\big{[}(G_{t}^{\top})^{2}+\epsilon\big{]}1_{m}$
     $\hat{V}_{t}=r_{t}s_{t}^{\top}/(1_{m}^{\top}r_{t})$
     $X_{t}=X_{t-1}-\eta_{t}\left(\mathrm{clip}\left(M_{t}/\sqrt{\hat{V}_{t}}\right)+\lambda X_{t-1}\right)$
  end for
Solving this problem results in a closed-form solution denoted by $r=V1_{m},s=V^{\top}1_{n}/r^{\top}1_{m}$. Subsequently, the iterative process of Adafactor optimizer can be outlined as in Algorithm  1, in which the decay parameter $\hat{\beta}_{2t}=\beta_{2}(1-\beta^{t-1}_{2})/(1-\beta^{t}_{2})$ is defined as equivalent to the bias correction step. In the model parameter update, we have $\mathrm{clip}(U)=U/\max(1,\mathrm{RMS}(U)/d)$ with $\mathrm{RMS}(.)$ refers to the root-mean-square function and $d$ is the threshold value, meaning that we cap the norm of the actual update rather than just the gradient. This technique aims to eliminate the larger-than-desired updates and also stabilize the training process with slow decay ($\beta_{2}=0.999$).
Briefly, Adafactor tracks the moving averages of the row and column sums of squared gradients throughout iterations, resulting in factored second-moment estimators $r_{t}$ and $s_{t}$. A normalized outer product $r_{t}s_{t}^{\top}/(1_{m}^{\top}r_{t})$ is then used to reconstruct a low-rank parameterization of second-order momentum. This technique is computationally efficient and scalable as it offers analytical formulations without the need for additional approximations. However, directly applying a non-negative decomposition approach like that often appears heuristic. It fails to provide any insights into the optimization dynamics and convergence guarantees when we reparameterize the second-order statistic using such a low-rank representation. Interestingly, we demonstrate that the iterative procedure in Adafactor-m optimizer can be discretized from an ordinary differential equation (ODE) as follows:
$\dot{X}$ $=-\frac{M}{\sqrt{rs^{\top}/1_{m}^{\top}r}},$ $\dot{M}$ $=\nabla f(X)-\alpha M,$ $\dot{r}$ $=(\nabla f(x))^{2}1_{n}-\alpha r,$ $\dot{s}$ $=(\nabla^{\top}f(x))^{2}1_{m}-\alpha s$ // Adafactor-m (ODE)
which solves a minimization problem with respect to the Hamiltonian function described by:
$H(X,M,r,s)=f(X)+\frac{1}{2}\sum_{i=1,j=1}^{m,n}\frac{M_{ij}^{2}\sqrt{\sum_{i=1}^{m}r_{i}}}{\sqrt{r_{i}s_{j}}}.$
Proposition 1.
A key property is that the function $H$ monotonically decreases along the ODE trajectory, that is, $\frac{\mathrm{d}}{\mathrm{d}t}H(X_{t},M_{t},r_{t},s_{t})\leq 0$.
A detailed proof is provided in Appendix A.2.
By formulating the update scheme as an ODE, we can elaborate on the dynamics of the optimization process, and more importantly, it facilitates theoretical insights into the stability and convergence properties. In the next sections, we will broaden this perspective to encompass a wider range of optimizers, including first-order and adaptive algorithms. By doing so, we aim to integrate various optimization techniques within a unified framework of Hamiltonian dynamics.
4.2 Factorized first-order momentum estimation
Factoring the second-order momentum is a straightforward yet intuitive idea to handle memory overhead. This approach offers an advantage in that it allows for the tolerance of information loss, as our focus is solely on the magnitude of the scaling parameter. On the other hand, an inherent drawback of most algorithms is the need to maintain a full momentum to avoid performance degradation. However, dealing with first-order information poses more challenging problems because we must consider both the magnitude and especially the update directions.
Factorization via rank-1 parameterization. We propose here a sublinear memory optimizer that factorizes the first-order momentum. The algorithm is specifically derived from an ordinary differential equation outlined as follows (recall that $X\in\mathbb{R}^{m\times n}$ is a matrix):
$\begin{split}\dot{X}&=-\nabla\phi\big{(}\beta\hat{u}1_{n}^{\top}+\nabla f(X)\big{)}-\nabla\psi\big{(}\beta 1_{m}\hat{v}^{\top}+\nabla f(X)\big{)}\\ \dot{u}&=\nabla f(X)1_{n}/n-\alpha u,~{}~{}~{}~{}~{}~{}~{}~{}~{}\hat{u}=u-\nabla f(X)1_{n}/n\\ \dot{v}&=\nabla^{\top}f(X)1_{m}/m-\alpha v,~{}~{}~{}~{}~{}~{}~{}~{}~{}\hat{v}=v-\nabla^{\top}f(X)1_{m}/m.\end{split}$ (2)
where $\alpha,\beta$ are two positive parameters; $\phi$ and $\psi$ are any convex functions, such that $\nabla\phi(\cdot),\nabla\psi(\cdot)$ are monotonic operators. The simplest option to consider is that $\phi(X)=\psi(X)=\left\lVert X\right\rVert_{2}^{2}/2$, then their gradients $\nabla\phi(X)=\nabla\psi(X)=X$. This system yields the Hamiltonian function:
$H(X,u,v)=f(X)+\frac{\beta n}{2}\left\lVert u\right\rVert_{2}^{2}+\frac{\beta m}{2}\left\lVert v\right\rVert_{2}^{2}.$
A detailed analysis of this function is provided in Appendix A.1.
Algorithm 2 signFSGD for matrix parameter, with factored first-order moments.
  Inputs: moment coefficients $\beta=0.9$, and regularization constant $\lambda$
  Initialization: weight parameters $X_{0}\in\mathbb{R}^{m\times n}$, initial moment factors $u_{0},v_{0}\leftarrow 0$
  for $t=1$ to $T$ do
     $G_{t}=\nabla f_{t}(X_{t-1})$
     $u_{t}=\beta u_{t-1}+(1-\beta)G_{t}1_{n}/n$
     $v_{t}=\beta v_{t-1}+(1-\beta)G_{t}^{\top}1_{m}/m$
     $\hat{u}_{t}=u_{t}-G_{t}1_{n}/n$
     $\hat{v}_{t}=v_{t}-G_{t}^{\top}1_{m}/m$
     $X_{t}=X_{t-1}-\eta_{t}\left(\mathrm{sign}\left(\beta\hat{u}_{t}1_{n}^{\top}+G_{t}\right)+\mathrm{sign}\left(\beta 1_{m}\hat{v}_{t}^{\top}+G_{t}\right)+\lambda X_{t-1}\right)$
  end for
(a) Validation accuracy on CIFAR datasets
(b) Validation accuracy on ImageNet1k
Figure 1: A comparison of optimizer performance on ResNet architectures. For signSGD, $m$ denotes the momentum coefficient. For signFSDG, “ablation" means the version without corrected terms, “fullhead" means the version using full momentum for the MLP head layer.
Back to our framework, a number of remarks are in order:
1) Compared with algorithms that maintain a full rank momentum matrix $M\in\mathbb{R}^{m\times n}$ in like equation 1, this algorithm employs two rank-one momentum vectors $u\in\mathbb{R}^{m}$ and $v\in\mathbb{R}^{n}$, which significantly reduces the memory cost from $O(mn)$ to $O(m+n)$. In the algorithm, we use the simple unit vectors $1_{m},1_{n}$ to first down project the full gradient $\nabla f(X)$ into the rank-one spaces of column means and row means, and exponentially accumulate these statistics across training steps. We then reconstruct the rank-one momentum $\hat{u},\hat{v}$ to update the model parameters. The $\hat{u}$ and $\hat{v}$ are centralized variants of $u$ and $v$.
2) Discretizing ODE equation 2 using the Euler method results in the following iterative scheme:
$u_{t}$ $=\hat{\beta}_{1t}u_{t-1}+(1-\hat{\beta}_{1t})\nabla f(X_{t})1_{n}/n,~{}~{}~{}~{}~{}~{}~{}~{}~{}\hat{u}_{t}=u_{t}-\nabla f(X_{t})1_{n}/n$
$v_{t}$ $=\hat{\beta}_{1t}v_{t-1}+(1-\hat{\beta}_{1t})\nabla^{\top}f(X_{t})1_{m}/m,~{}~{}~{}~{}~{}~{}~{}~{}~{}\hat{v}_{t}=v_{t}-\nabla^{\top}f(X_{t})1_{m}/m$
$X_{t}$ $=X_{t-1}-\eta_{t}\bigg{[}\nabla\phi\left(\hat{\beta}_{1t}\hat{u}_{t}1_{n}^{\top}+\nabla f(X_{t})\right)+\nabla\psi\left(\hat{\beta}_{1t}1_{m}\hat{v}_{t}^{\top}+\nabla f(X_{t})\right)\bigg{]}$
In discrete-time analysis, we centralize the row means and column means statistics $u_{t},v_{t}$ by corrected terms $\nabla f(X_{t})1_{n}/n$ and $\nabla^{\top}f(X_{t})1_{m}/m$, respectively. These corrected terms are essential as they can guarantee the updates inside $\nabla\phi(.)$ and $\nabla\psi(.)$ will converge to the true first moment $\mathbb{E}[\nabla f(X_{t})]$. Specifically, by unfolding the factored moment $u_{t}$ (similarly to $v_{t}$), we can get an accumulator of column-wise mean of gradients in the form:
$u_{t}=\sum_{i=1}^{t}(1-\hat{\beta}_{1i})\prod_{j=i+1}^{t}\hat{\beta}_{1j}\nabla f(X_{i})1_{n}.$
By the assumption that the gradient distribution is stationary, taking expectations gives us:
$\mathbb{E}[u_{t}]=\sum_{i=1}^{t}(1-\hat{\beta}_{1i})\prod_{j=i+1}^{t}\hat{\beta}_{1j}\mathbb{E}[\nabla f(X_{i})1_{n}]=\sum_{i=1}^{t}(1-\hat{\beta}_{1i})\prod_{j=i+1}^{t}\hat{\beta}_{1j}\mathbb{E}[\nabla f(X_{t})1_{n}].$
Futhermore, we can prove by induction that $\sum_{i=1}^{t}(1-\hat{\beta}_{1i})\prod_{j=i+1}^{t}\hat{\beta}_{1j}=1$, leading to $\mathbb{E}[u_{t}]=\mathbb{E}[\nabla f(X_{t})1_{n}]$ and hence the expected value of centralized moment factor $\mathbb{E}[\hat{u}_{t}]$ will approximate the true first-order moment $\mathbb{E}[\nabla f(X_{t})]$.
3) When the momentum coefficient is deactivated ($\beta_{1}=0$), the update terms inside $\nabla\phi(.)$ and $\nabla\psi(.)$ degenerate to $\nabla f(X_{t})$. In this scenario, the algorithm simplifies to standard gradient descent:
$\dot{X}=-\nabla\phi(\nabla f(X))-\nabla\psi(\nabla f(X)).$
[…]
4.3 Fully factorized momentum estimations
Figure 2: Histograms illustrating the gradients of MLP head layers in ResNet50 (top) and ResNet101 (bottom) trained on CIFAR10 and CIFAR100, respectively.
Inadequacy of first-order factorization. For ResNet models, mini-batch gradients are typically quite small and well concentrated around zero mean as shown in Figure 2. Consequently, accumulating the momentum through column means and row means of the gradient information will significantly flatten the magnitude. This issue could potentially reduce the impact of momentum on the optimization updates, especially with large-scale settings such as ImageNet and modern deep learning architectures.
In Figure 1(b), we present the performances of signSGD and signFSGD when training ResNet50 from scratch on ImageNet1k. Notably, our method does not yield substantial enhancements compared to signSGD without momentum as observed on CIFAR10, CIFAR100 datasets. However, when employing a full momentum instead of the factorized one for the MLP head layer, our signFSGD can perform comparably to signSGD with momentum. We witness the same behavior when applying the algorithm to larger ResNet architectures.
This observation reinforces our previous remark regarding the moderate impact of our factorized first-order momentum on scale models. Although a few minor customizations, such as omitting the factorization of the MLP head layer like above, could overcome the issue on ResNets, our initial expectation leaned towards a more general-purpose optimization algorithm. We specifically aim to extend the scalability of our algorithms to encompass modern convolution-free architectures, such as Vision Transformers.
Incorporating second-order information. A natural idea to deal with the magnitude issue of factorized first-order momentum is to normalize them by factored second moments. Indeed, by integrating our factorization for the first moment estimator into Adafactor, we can establish a unified method where both momentum estimators are fully factorized. In principle, we aim to conceptualize our algorithms within Hamiltonian frameworks, but deriving an objective function for such a simplistic integration might be challenging. We instead come up with a novel algorithm characterized by the following ODE:
$\dot{X}$ $=-\frac{1}{2}\bigg{(}\frac{u1_{n}^{\top}-\nabla f(X)1_{n}1_{n}^{\top}/n}{\sqrt{r1_{n}^{\top}}}+\frac{1_{m}v^{\top}-1_{m}1_{m}^{\top}\nabla f(X)/m}{\sqrt{1_{m}s^{\top}}}\bigg{)}-\frac{\nabla f(X)}{\sqrt{rs^{\top}/1_{m}^{\top}r}}$
$\dot{u}$ $=\nabla f(X)1_{n}/n-\alpha u$
$\dot{v}$ $=\nabla^{\top}f(X)1_{m}/m-\alpha v$
$\dot{r}$ $=(\nabla f(X))^{2}1_{n}-\alpha r$
$\dot{s}$ $=(\nabla^{\top}f(X))^{2}1_{m}-\alpha s$ // H-Fac (ODE)
which yields the following Hamiltonian function:
$H(X,u,v,r,s):=f(X)+\frac{n}{4}\sum_{i=1}^{m}\frac{u_{i}^{2}}{\sqrt{r_{i}}}+\frac{m}{4}\sum_{j=1}^{n}\frac{v_{j}^{2}}{\sqrt{s_{j}}}.$
Proposition 2.
The function $H$ monotonically decreases along the ODE trajectory.
A detailed proof is provided in Appendix A.3.
Algorithm 3 H-Fac for matrix parameters, with both first and second-order moments factorized.
  Inputs: moment decay coefficients $\beta_{1},\beta_{2}$, smoothing term $\epsilon$, and regularization constant $\lambda$
  Initialization: weight parameters $X_{0}\in\mathbb{R}^{m\times n}$, initial factored moments $u_{0},v_{0},r_{0},s_{0}\leftarrow 0$
  for $t=1$ to $T$ do
     $G_{t}=\nabla f_{t}(X_{t-1})$
     $u_{t}=\hat{\beta}_{1t}u_{t-1}+(1-\hat{\beta}_{1t})G_{t}1_{n}/n$
     $v_{t}=\hat{\beta}_{1t}v_{t-1}+(1-\hat{\beta}_{1t})G_{t}^{\top}1_{m}/m$
     $r_{t}=\hat{\beta}_{2t}r_{t-1}+(1-\hat{\beta}_{2t})\big{[}(G_{t})^{2}+\epsilon\big{]}1_{n}$
     $s_{t}=\hat{\beta}_{2t}s_{t-1}+(1-\hat{\beta}_{2t})\big{[}(G_{t}^{\top})^{2}+\epsilon\big{]}1_{m}$
     $\hat{V}_{t}=r_{t}s_{t}^{\top}/(1_{m}^{\top}r_{t})$
     $\phi_{term}=\hat{\beta}_{1t}\big{(}u_{t}1_{n}^{\top}-G_{t}1_{n}1_{n}^{\top}/n\big{)}/\sqrt{{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}r_{t}1_{n}^{\top}/n}}}$
     $\psi_{term}=\hat{\beta}_{1t}\big{(}1_{m}v_{t}^{\top}-1_{m}1_{m}^{\top}G_{t}/m\big{)}/\sqrt{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1_{m}s_{t}^{\top}/m}}$
     $X_{t}=X_{t-1}-\eta_{t}\left(0.5(\phi_{term}+\psi_{term})+\mathrm{clip}\left(G_{t}/\sqrt{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\hat{V}_{t}}}\right)+\lambda X_{t-1}\right)$
  end for
The discrete-time equivalent is presented in Algorithm 3. In fact, this algorithm is inspired by a variant of AdamW optimizer outlined in Appendix C. Our model parameters update can be seen as an accumulator of normalized gradients, where both the momentum and the current gradient are normalized by their corresponding cumulative second-moment information. More specifically, our update includes a combination of three key elements: (1) the normalized momentum factorization $0.5*(\phi_{term}+\psi_{term})$, in which the factorized first-order moment in section 4.2 is normalized by the row means and the column means of second-moment estimators; (2) a clipping of normalized gradient $\mathrm{clip}\big{(}G_{t}/\sqrt{\hat{V}_{t}}\big{)}$ which is inherited from Adafactor update; and (3) a decouple weight decay $\lambda X_{t-1}$ for enhancing the generalization performance of adaptive optimizers [24].
5 Experiments
In this section, we conduct several experiments to demonstrate the effectiveness of our H-Fac algorithm. Unlike the earlier proposal of signFSGD, we do not customize the algorithm for any specific layers, but perform factorization on the entire architectures.
Experimental Setup
We evaluated the optimization algorithms described in this paper mainly on the image classification task, specifically training ResNet50 and Vision Transformers from scratch on ImageNet1k, following previous works [11, 16]. The images are pre-processed by Inception-style cropping [33] and random horizontal. We train ResNet50 for 90 epochs, using a batch size of 1024, with cosine learning rate decay scheduler. For Vision Transformers (ViTs), we train them for 300 epochs, using a batch size of 4096, with a learning rate schedule of 10,000 steps warmup followed by linear decay. We also adopted strong data augmentations, including RandAugment( 2,15) [7] and mixup (0.5) [39], to boost ViTs performances. We opted for recommended configurations from prior research for hyperparameters such as learning rate ($lr$) and weight decay ($\lambda$), dropout ($dr$):
  • $lr=3e-4$, $\lambda=1.0$ in sign-based optimizers, $lr=1e-3$, $\lambda=0.1$ in adaptive algorithms when training ResNet50 on ImageNet.
  • $lr=3e-3$, $\lambda=0.1$, $dr=0$ in adaptive algorithms for training ViT-B/32, ViT-S/16 on ImageNet with strong augmentations.
Figure 3: Top-1 Accuracy of optimizers in training ResNet50, ViT-B/32, and ViT-S/16 from scratch on the ImageNet1k. For H-Fac, “ablation" means the version without corrected terms.
Table 1: ImageNet1k top-1 accuracy results for various optimizers on different models.
Optimizers signSGD signFSGD signFSGD AdamW Adafactor H-Fac
$m=0.9$ $m=0.0$fullhead $m=0.9$ $m=0.0$
ResNet50 74.47 72.36 72.47 74.45 75.66 75.85 75.79 75.90
ViT-B/32 72.20 72.31 71.36 71.87
ViT-S/16 78.16 77.81 76.74 77.20
5.1 Empirical Results
The results are shown in Figure 3 and Table 1. On ResNet50, we can notice consistent patterns with those in Figure 1(b). While our algorithm, H-Fac, can achieve similar memory efficiency as Adafactor without momentum, it demonstrates more stable training and delivers highly competitive performance compared to Adafactor with momentum. A performance degradation is also clearly observed in the ablation version, consolidating the significance of the corrected terms in our algorithms. On ViT models, although our H-Fac can catch up AdamW and Adafactor with momentum, there are noticeable performance gaps between optimizers, especially during the early training stages. A positive sign is that the factorized momentum in our algorithm can actually accelerate the optimization process, as evidenced by significant improvements compared to Adafactor without momentum. It should be noted that our signFSGD algorithm does not perform well on ViT architectures, even when full momentum is applied to the head layer. This result is anticipated according to the inadequacies discussed in Section 4.3.
We can elaborate on these behaviors by exploring the trainability of ResNets and ViTs. Convolution-free architectures like ViTs and MLP-Mixer are theoretically more difficult to train, as highlighted by analyses of the neural tangent kernel (NTK) condition number [5, 38]. Additionally, Park et al. [28] further discovered through Hessian spectrum analysis that ResNets possess a locally convex loss landscape, while ViTs exhibit many directions corresponding to negative Hessian eigenvalues. These aspects imply the need for better-adapted momentum in optimizing convolution-free networks, and partially explain the varied effects of our factorized momentum on the architectures. It also makes sense to expect that leveraging new architectural innovations or methods for smoothing the loss landscape could improve our optimization algorithms.
6 Limitation and Discussion
We conducted a few large-scale experiments on language models (LLMs) but unfortunately, we found our factorized optimizers might not be very effective. In the sense that our H-Fac optimizer only gains negligible improvement compared to Adafactor without momentum. As the model size increases considerably, the rank-1 parameterization in our factorized momentum only offers very marginal contributions, which is insufficient to accelerate the optimization process as effectively as proper momentum. This suggests that developing algorithms with rank-$k$ approximation, based on the same principles of Hamiltonian dynamics, can hold significant potential for future works.
In conclusion, our work in this paper represents one of the first attempts to unify and devise new adaptive optimization methods that are theoretically grounded in Hamiltonian frameworks. It would be exciting to see future research pursuing this direction to enhance understanding of existing algorithms and propose more efficient ones. For practical use, our current optimizers show promise in applications where ResNets are advantageous because of their efficient training nature. For example, we can adapt our algorithms to the federated learning optimization problem, so that the factored moment estimators can be efficiently communicated to accelerate the convergence. Other applications would also be of interest.
Appendix A Hamiltonian function analyses
A.1 Factorized first-order cases
Recall our general framework for factorized first-order methods, which is characterized by the following ODE:
$\begin{split}\dot{X}&=-\nabla\phi\big{(}\beta\hat{u}1_{n}^{\top}+\nabla f(X)\big{)}-\nabla\psi\big{(}\beta 1_{m}\hat{v}^{\top}+\nabla f(X)\big{)}\\ \dot{u}&=\nabla f(X)1_{n}/n-\alpha u,~{}~{}~{}~{}~{}~{}~{}~{}~{}\hat{u}=u-\nabla f(X)1_{n}/n\\ \dot{v}&=\nabla^{\top}f(X)1_{m}/m-\alpha v,~{}~{}~{}~{}~{}~{}~{}~{}~{}\hat{v}=v-\nabla^{\top}f(X)1_{m}/m.\end{split}$
For a canonical example with $\phi(X)=\psi(X)=\left\lVert X\right\rVert_{2}^{2}/2$, and hence $\nabla\phi(X)=\nabla\psi(X)=X$. We will show that the following Hamiltonian function:
$H(X,u,v)=f(X)+\frac{\beta n}{2}\left\lVert u\right\rVert_{2}^{2}+\frac{\beta m}{2}\left\lVert v\right\rVert_{2}^{2},$
monotonically decreases along the ODE trajectory above. Denote $G=\nabla f(X)$, and note that we can explicitly write the derivative of the model parameter:
$\dot{X}=-\left(\beta u1_{n}^{\top}+G-\frac{\beta}{n}G1_{n}1_{n}^{\top}\right)-\left(\beta 1_{m}v^{\top}+G-\frac{\beta}{m}1_{m}1_{m}^{\top}G\right),$
then we can shorten the derivative of the function $H$ as:
$\frac{\mathrm{d}}{\mathrm{d}t}H$ $=\text{trace}(G^{\top}\dot{X})+\frac{\beta}{n}u^{\top}\dot{u}+\frac{\beta}{m}v^{\top}\dot{v}$
$=-\text{trace}\left(G^{\top}\left(G-\frac{\beta}{n}G1_{n}1_{n}^{\top}\right)\right)-\text{trace}\left(G^{\top}\left(G-\frac{\beta}{m}1_{m}1_{m}^{\top}G\right)\right)-\frac{\alpha\beta}{n}\left\lVert u\right\rVert_{2}^{2}-\frac{\alpha\beta}{m}\left\lVert v\right\rVert_{2}^{2}$
$=-2\left\lVert G\right\rVert_{F}^{2}+\frac{\beta}{n}\left\lVert G1_{n}\right\rVert_{2}^{2}+\frac{\beta}{m}\left\lVert G^{\top}1_{m}\right\rVert_{2}^{2}-\frac{\alpha\beta}{n}\left\lVert u\right\rVert_{2}^{2}-\frac{\alpha\beta}{m}\left\lVert v\right\rVert_{2}^{2},$
in which .$\left\lVert.\right\rVert_{F}$ is the Frobenius norm. Since $\beta\in(0,1)$, applying C-S inequality, we have:
$\frac{\mathrm{d}}{\mathrm{d}t}H\leq-2\left\lVert G\right\rVert_{F}^{2}+\frac{1}{n}\left\lVert G1_{n}\right\rVert_{2}^{2}+\frac{1}{m}\left\lVert G^{\top}1_{m}\right\rVert_{2}^{2}\leq 0.$
A.2 Adafactor
Recall the ODE of Adafactor with momentum:
$\dot{X}$ $=-\frac{M}{\sqrt{rs^{\top}/1_{m}^{\top}r}},$ $\dot{M}$ $=\nabla f(X)-\alpha M,$ $\dot{r}$ $=(\nabla f(X))^{2}1_{n}-\alpha r,$ $\dot{s}$ $=(\nabla^{\top}f(X))^{2}1_{m}-\alpha s.$
We will show that the Hamiltonian function described by:
$H(X,M,r,s)=f(X)+\underbrace{\frac{1}{2}\sum_{i=1,j=1}^{m,n}\frac{M_{ij}^{2}\sqrt{\sum_{i=1}^{m}r_{i}}}{\sqrt{r_{i}s_{j}}}}_{\Phi(M,r,s)},$
is monotonically decreasing along the ODE trajectory.
Denote $G=\nabla f(X)$, we have the derivative of $H$ can be expressed by:
$\frac{\mathrm{d}}{\mathrm{d}t}H=\text{trace}(G^{\top}\dot{X})+\text{trace}\left((\nabla_{M}\Phi)^{\top}\dot{M}\right)+(\nabla_{r}\Phi)^{\top}\dot{r}+(\nabla_{s}\Phi)^{\top}\dot{s}$
Analyzing each element specifically, we get:
$\text{trace}(G^{\top}\dot{X})$ $=-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{trace}\left(G^{\top}\frac{M\sqrt{1_{m}^{\top}r}}{\sqrt{rs^{\top}}}\right)},$
$\text{trace}\left((\nabla_{M}\Phi)^{\top}\dot{M}\right)$ $=\text{trace}\left((G-\alpha M)^{\top}\frac{M\sqrt{1_{m}^{\top}r}}{\sqrt{rs^{\top}}}\right)={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{trace}\left(G^{\top}\frac{M\sqrt{1_{m}^{\top}r}}{\sqrt{rs^{\top}}}\right)}-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\alpha\text{trace}\left(M^{\top}\frac{M\sqrt{1_{m}^{\top}r}}{\sqrt{rs^{\top}}}\right)},$
$(\nabla_{r}\Phi)^{\top}\dot{r}$ $=\sum_{i=1}^{m}(\partial_{r_{i}}\Phi)(1_{n}^{\top}G_{i}^{2}-\alpha r_{i})$ // $G_{i}$ is the $i$’th row of $G$
$=\sum_{i=1}^{m}\left(\frac{1}{2}\sum_{j=1}^{n}\frac{M_{ij}^{2}}{\sqrt{s_{j}}}\right)\frac{-\sum_{k\neq i}r_{k}}{2\sqrt{r_{i}^{3/2}}\sqrt{\sum_{i}r_{i}}}(1_{n}^{\top}G_{i}^{2}-\alpha r_{i})$
$\leq\sum_{i=1}^{m}\left(\frac{\alpha}{4}\sum_{j=1}^{n}\frac{M_{ij}^{2}}{\sqrt{s_{j}}}\right)\frac{\sum_{k\neq i}r_{k}}{\sqrt{r_{i}}\sqrt{\sum_{i}r_{i}}}$ // the multiplication by $1_{n}^{\top}G_{i}^{2}$ is $\leq 0$
$\leq\sum_{i=1}^{m}\left(\frac{\alpha}{4}\sum_{j=1}^{n}\frac{M_{ij}^{2}}{\sqrt{s_{j}}}\right)\frac{\sqrt{\sum_{i}r_{i}}}{\sqrt{r_{i}}}$ // $\sum_{k\neq i}r_{k}\leq\sum_{i}r_{i}$
$=\sum_{i=1}^{m}\left(\frac{\alpha}{4}\sum_{j=1}^{n}\frac{M_{ij}^{2}\sqrt{\sum_{i}r_{i}}}{\sqrt{r_{i}s_{j}}}\right)$
$={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\frac{\alpha}{4}\text{trace}\left(M^{\top}\frac{M\sqrt{1_{m}^{\top}r}}{\sqrt{rs^{\top}}}\right)},$
$(\nabla_{s}\Phi)^{\top}\dot{s}$ $=\sum_{j=1}^{n}(\partial_{s_{j}}\Phi)(1_{m}^{\top}G_{:j}^{2}-\alpha s_{j})$ // $G_{:j}$ is the $j$’th column of $G$
$=\sum_{j=1}^{n}\left(\frac{1}{2}\sum_{i=1}^{m}\frac{M_{ij}^{2}\sqrt{\sum_{i}r_{i}}}{\sqrt{r_{i}}}\right)\frac{-1}{2\sqrt{s_{j}^{3/2}}}(1_{m}^{\top}G_{:j}^{2}-\alpha s_{j})$
$\leq\sum_{j=1}^{n}\left(\frac{\alpha}{4}\sum_{i=1}^{m}\frac{M_{ij}^{2}\sqrt{\sum_{i}r_{i}}}{\sqrt{r_{i}s_{j}}}\right)$ // the multiplication by $1_{m}^{\top}G_{:j}^{2}$ is $\leq 0$
$={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\frac{\alpha}{4}\text{trace}\left(M^{\top}\frac{M\sqrt{1_{m}^{\top}r}}{\sqrt{rs^{\top}}}\right)}$
Canceling out similar quantities, we can rewrite the Hamiltonian derivative as follows:
$\frac{\mathrm{d}}{\mathrm{d}t}H\leq-\frac{\alpha}{2}\text{trace}\left(M^{\top}\frac{M\sqrt{1_{m}^{\top}r}}{\sqrt{rs^{\top}}}\right)=-\frac{\alpha}{2}\sum_{i=1}^{m}\sum_{j=1}^{n}\frac{M_{ij}^{2}\sqrt{\sum_{i}r_{i}}}{\sqrt{r_{i}s_{j}}}\leq 0.$
A.3 H-Fac
In this part, we will prove that the following ODE trajectory:
$\dot{X}$ $=-\frac{1}{2}\bigg{(}\frac{u1_{n}^{\top}-\nabla f(X)1_{n}1_{n}^{\top}/n}{\sqrt{r1_{n}^{\top}}}+\frac{1_{m}v^{\top}-1_{m}1_{m}^{\top}\nabla f(X)/m}{\sqrt{1_{m}s^{\top}}}\bigg{)}-\frac{\nabla f(X)}{\sqrt{rs^{\top}/1_{m}^{\top}r}}$
$\dot{u}$ $=\nabla f(X)1_{n}/n-\alpha u$
$\dot{v}$ $=\nabla^{\top}f(X)1_{m}/m-\alpha v$
$\dot{r}$ $=(\nabla f(X))^{2}1_{n}-\alpha r$
$\dot{s}$ $=(\nabla^{\top}f(X))^{2}1_{m}-\alpha s$
descends the Hamiltonian function defined by:
$H(X,u,v,r,s):=f(X)+\underbrace{\frac{n}{4}\sum_{i=1}^{m}\frac{u_{i}^{2}}{\sqrt{r_{i}}}+\frac{m}{4}\sum_{j=1}^{n}\frac{v_{j}^{2}}{\sqrt{s_{j}}}}_{\Phi(u,v,r,s)}.$
Denote $G=\nabla f(X)$, we have:
$\frac{\mathrm{d}}{\mathrm{d}t}H=\text{trace}(G^{\top}\dot{X})+(\nabla_{u}\Phi)^{\top}\dot{u}+(\nabla_{v}\Phi)^{\top}\dot{v}+(\nabla_{r}\Phi)^{\top}\dot{r}+(\nabla_{s}\Phi)^{\top}\dot{s}$
Similar to the previous part, we can calculate each element specifically as follows:
$\text{trace}(G^{\top}\dot{X})$ $=-\frac{1}{2}\text{trace}\left(G^{\top}\frac{u1_{n}^{\top}}{\sqrt{r1_{n}^{\top}}}\right)-\frac{1}{2}\text{trace}\left(G^{\top}\frac{1_{m}^{\top}v}{\sqrt{1_{m}s^{\top}}}\right)$
$\quad+\text{trace}\left(G^{\top}\frac{G1_{n}1_{n}^{\top}}{2n\sqrt{r1_{n}^{\top}}}\right)+\text{trace}\left(G^{\top}\frac{1_{m}1_{m}^{\top}G}{2m\sqrt{1_{m}s^{\top}}}\right)-\text{trace}\left(G^{\top}\frac{G\sqrt{1_{m}^{\top}r}}{\sqrt{rs^{\top}}}\right)$
$(\nabla_{u}\Phi)^{\top}\dot{u}$ $=\frac{n}{2}\frac{u^{\top}}{\sqrt{r^{\top}}}(G1_{n}/n-\alpha u)=\frac{1}{2}\frac{u^{\top}}{\sqrt{r^{\top}}}G1_{n}-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\frac{n\alpha}{2}\frac{u^{\top}}{\sqrt{r^{\top}}}u}$
$(\nabla_{v}\Phi)^{\top}\dot{v}$ $=\frac{m}{2}\frac{v^{\top}}{\sqrt{s^{\top}}}(G^{\top}1_{m}/m-\alpha v)=\frac{1}{2}\frac{v^{\top}}{\sqrt{s^{\top}}}G^{\top}1_{m}-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\frac{m\alpha}{2}\frac{v^{\top}}{\sqrt{s^{\top}}}v}$
$(\nabla_{r}\Phi)^{\top}\dot{r}$ $=\left(-\frac{n}{8}\frac{u^{2}}{\sqrt{r^{3/2}}}\right)^{\top}(G^{2}1_{n}-\alpha r)=\frac{n\alpha}{8}\frac{u^{2}}{\sqrt{r}}-\left(\frac{n}{8}\frac{u^{2}}{\sqrt{r^{3/2}}}\right)^{\top}G^{2}1_{n}\leq{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\frac{n\alpha}{8}\frac{u^{2}}{\sqrt{r}}}$
$(\nabla_{s}\Phi)^{\top}\dot{s}$ $=\left(-\frac{m}{8}\frac{v^{2}}{\sqrt{s^{3/2}}}\right)^{\top}((G^{2})^{\top}1_{m}-\alpha s)=\frac{m\alpha}{8}\frac{v^{2}}{\sqrt{s}}-\left(\frac{m}{8}\frac{v^{2}}{\sqrt{s^{3/2}}}\right)^{\top}(G^{2})^{\top}1_{m}\leq{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\frac{m\alpha}{8}\frac{v^{2}}{\sqrt{s}}}$
Canceling out crossing terms, with a note that:
$\text{trace}\left(G^{\top}\frac{u1_{n}^{\top}}{\sqrt{r1_{n}^{\top}}}\right)$ $=\frac{u^{\top}}{\sqrt{r^{\top}}}G1_{n},~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{trace}\left(G^{\top}\frac{1_{m}^{\top}v}{\sqrt{1_{m}s^{\top}}}\right)=\frac{v^{\top}}{\sqrt{s^{\top}}}G^{\top}1_{m},$
$\frac{u^{\top}}{\sqrt{r^{\top}}}u$ $=\frac{u^{2}}{\sqrt{r}},~{}~{}~{}~{}~{}~{}~{}~{}~{}\frac{v^{\top}}{\sqrt{s^{\top}}}v=\frac{v^{2}}{\sqrt{s}},$
we can rewrite the Hamiltonian derivative as follows:
$\frac{\mathrm{d}}{\mathrm{d}t}H$ $\leq\text{trace}\left(G^{\top}\frac{G1_{n}1_{n}^{\top}}{2n\sqrt{r1_{n}^{\top}}}\right)+\text{trace}\left(G^{\top}\frac{1_{m}1_{m}^{\top}G}{2m\sqrt{1_{m}s^{\top}}}\right)-\text{trace}\left(G^{\top}\frac{G\sqrt{1_{m}^{\top}r}}{\sqrt{rs^{\top}}}\right)$
$=\text{trace}\left(\frac{1}{n}(G1_{n})^{2}\frac{1}{2\sqrt{r^{\top}}}\right)+\text{trace}\left(\frac{1}{m}(G^{T}1_{m})^{2}\frac{1}{2\sqrt{s^{\top}}}\right)-\text{trace}\left(G^{2}\frac{\sqrt{1_{m}^{\top}r}}{\sqrt{sr^{\top}}}\right).$
Applying C-S inequality gives us $\frac{1}{n}(G1_{n})^{2}\leq G^{2}1_{n}$ and $\frac{1}{m}(G^{T}1_{m})^{2}\leq(G^{T})^{2}1_{m}$, therefore:
$\frac{\mathrm{d}}{\mathrm{d}t}H$ $\leq\text{trace}\left(G^{2}1_{n}\frac{1}{2\sqrt{r^{\top}}}\right)+\text{trace}\left((G^{T})^{2}1_{m}\frac{1}{2\sqrt{s^{\top}}}\right)-\text{trace}\left(G^{2}\frac{\sqrt{1_{m}^{\top}r}}{\sqrt{sr^{\top}}}\right),$
$=\frac{1}{2}\text{trace}\bigg{[}G^{2}\bigg{(}\underbrace{1_{n}\frac{1}{\sqrt{r^{\top}}}+\frac{1}{\sqrt{s}}1_{m}^{\top}-2\frac{\sqrt{1_{m}^{\top}r}}{\sqrt{sr^{\top}}}}_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}Q}}\bigg{)}\bigg{]}.$
By zero initialization, we note that the moving averages of row sums and column sums are symmetric, in other words, $1_{m}^{\top}r$ and $1_{n}^{\top}s$ are both represent the moving average of the sum of all squared gradient entries. Denote this quantity by $\mathcal{S}$, we have $\mathcal{S}=1_{m}^{\top}r=1_{n}^{\top}s\geq(r_{i}+s_{j})/2$ for all $i,j$, then consider each element of matrix $Q$:
$\frac{1}{\sqrt{r_{i}}}+\frac{1}{\sqrt{s_{j}}}-2\frac{\sqrt{\mathcal{S}}}{\sqrt{r_{i}s_{j}}}\leq\frac{1}{\sqrt{r_{i}}}+\frac{1}{\sqrt{s_{j}}}-\frac{\sqrt{2(r_{i}+s_{j})}}{\sqrt{r_{i}s_{j}}}\leq 0,$
by C-S inequality. Therefore $Q$ is a negative matrix, and as a result, $\frac{\mathrm{d}}{\mathrm{d}t}H(x_{t},u_{t},v_{t},r_{t},s_{t})\leq 0$. ∎
Appendix B Factorized Lion optimizer
Recently, a new optimization named Lion (Evolved Sign Momentum) [6] was discovered by an evolutionary search algorithm applied to a symbolically represented program space. Lion has been shown to achieve at least comparable performance to AdamW on a wide range of tasks while reducing memory cost and training time. Notably, Lion can be formulated as an iterative update procedure:
$M_{t}$ $=\beta_{2}M_{t-1}+(1-\beta_{2})\nabla f(X_{t-1})$
$X_{t}$ $=X_{t-1}-\eta_{t}\left(\text{sign}(\beta_{1}M_{t-1}+(1-\beta_{1})\nabla f(X_{t-1}))+\lambda X_{t-1}\right)$ // Lion
[…]
Algorithm 4 Lionfactor for matrix parameter, with factored first-order moments.
  Inputs: double-moment coefficients $\beta_{1}=0.9,\beta_{2}=0.99$, and regularization constant $\lambda$
  Initialization: weight parameters $X_{0}\in\mathbb{R}^{m\times n}$, initial moment factors $u_{0},v_{0}\leftarrow 0$
  for $t=1$ to $T$ do
     $G_{t}=\nabla f_{t}(X_{t-1})$
     update model parrameters
     $\hat{u}_{t}=\beta_{1}u_{t-1}+(1-\beta_{1})G_{t}1_{n}/n-G_{t}1_{n}/n$
     $\hat{v}_{t}=\beta_{1}v_{t-1}+(1-\beta_{1})G_{t}^{\top}1_{m}/m-G_{t}^{\top}1_{m}/m$
     $X_{t}=X_{t-1}-\eta_{t}\left(\mathrm{sign}\left(\hat{u}_{t}1_{n}^{\top}+G_{t}\right)+\mathrm{sign}\left(1_{m}\hat{v}_{t}^{\top}+G_{t}\right)+\lambda X_{t-1}\right)$
     update exponential moment averages
     $u_{t}=\beta_{2}u_{t-1}+(1-\beta_{2})G_{t}1_{n}/n$
     $v_{t}=\beta_{2}v_{t-1}+(1-\beta_{2})G_{t}^{\top}1_{m}/m$
  end for
Figure 4: Performance of sign-based optimizers on ResNet architectures. “fullhead" means the version using full momentum for the MLP head layer.
Appendix C Pseudocode for Adam and a variant
Algorithm 5 AdamW
  Inputs: moment coefficients $\beta_{1},\beta_{2}$, weight decay constant $\lambda$, and smoothing term $\epsilon$.
  Initialization: weight parameter $X_{0}\in\mathbb{R}^{m\times n}$, initial moments $M_{0},V_{0}\leftarrow 0$
  for $t=1$ to $T$ do
     $G_{t}=\nabla f_{t}(X_{t-1})$
     $M_{t}=\hat{\beta}_{1t}M_{t-1}+(1-\hat{\beta}_{1t})G_{t}$
     $V_{t}=\hat{\beta}_{2t}V_{t-1}+(1-\hat{\beta}_{2t})G_{t}^{2}$
     $X_{t}=X_{t-1}-\eta_{t}\left(M_{t}/\left(\sqrt{V_{t}}+\epsilon\right)+\lambda X_{t-1}\right)$
  end for
Algorithm 6 AdamW variant
  Inputs: moment coefficients $\beta_{1},\beta_{2}$, weight decay constant $\lambda$, and smoothing term $\epsilon$.
  Initialization: weight parameter $X_{0}\in\mathbb{R}^{m\times n}$, initial moments $M_{0},V_{0}\leftarrow 0$
  for $t=1$ to $T$ do
     $G_{t}=\nabla f_{t}(X_{t-1})$
     $V_{t}=\hat{\beta}_{2t}V_{t-1}+(1-\hat{\beta}_{2t})G_{t}^{2}$
     $M_{t}=\hat{\beta}_{1t}M_{t-1}+(1-\hat{\beta}_{1t})G_{t}/\left(\sqrt{V_{t}}+\epsilon\right)$
     $X_{t}=X_{t-1}-\eta_{t}(M_{t}+\lambda X_{t-1})$
  end for
We reproduce the pseudocode for the Adam optimizer and its variant. Basically, Adam adopts exponential moving averages to accumulate the first ($M_{t}$) and second-order momentum ($V_{t}$) through training steps, and then uses the signal-to-noise ratio $M_{t}/\sqrt{V_{t}}$ to update the model parameters. While the variant at each iteration, will utilize the second-moment estimator to normalize the current gradient, then exponentially accumulate this normalized value for updating parameters. Our H-Fac shares the same principles as this variant.